# Non-Iterative Coordination of Interconnected Power Grids via Dimension-Decomposition-Based Flexibility Aggregation

### v2

@@ IV-B Physical Counterpart of the Flexibility Aggregation Model
The flexibility aggregation model derived through geometric projection calculations can be interpreted in terms of physical significance, as the parameters of the aggregated flexibility region of the RPG have a clear physical meaning.
For a single tie-line, its flexibility region across all time slots can be represented by an equivalent generator model, constrained by both power and ramp rate constraints, corresponding to the temporal coupling constraints in (18).
For a single time slice, the output power and bus voltage angles of all equivalent generators are interconnected through the constraints in the ellipsoidal expression form, corresponding to the single time slice coupling constraints in (21).
Therefore, after flexibility aggregation, the entire RPG can be equivalent to a group of generators that interconnected through coupling constraints, as shown in the schematic diagram Fig. 7.
Figure 7: (a) The original regional power grid; (b) Physical counterpart: equivalent generators model via flexibility aggregation.
IV-C Aggregated Operational Cost Function
To facilitate coordinated economic dispatch of interconnected RPGs, the analytical aggregated cost function of RPG is derived with the piecewise linear fitting [28]. The total operational cost of an RPG is the sum of the operational costs of each generation unit. From an external system perspective, the operational cost of an RPG, denoted as $\kappa_{t}$, is a function of the power flow through all tie-lines, denoted as $\bm{p}^{\text{tie}}_{t}$. To model this aggregated cost function, uniform sampling is applied across each dimension of the feasible domain, yielding a total of $J$ samples of tie-line power flows and their corresponding operational costs. The operating point and the corresponding cost of the $j$-th sample are denoted by $\bm{p}^{\text{tie}}_{t(j)}$ and $\kappa_{t(j)}$, respectively. By solving the following optimal power flow problem, the minimum operational cost $\kappa_{t(j)}$ for the $j$-th operating point $\bm{p}^{\text{tie}}_{t(j)}$ can be obtained:
@@ V Coordinated Dispatch of Interconnected RPGs Based on the Aggregated Model
With the flexibility aggregated models (28) and the aggregated cost function (30) of the RPGs, the coordinated dispatch of multiple interconnected RPGs can be formulated as follows
@@ VI-A Simulation Setup
Numerical tests are conducted on two cases. A simple IEEE 118 system with a DC tie-line and an AC tie-line is used to demonstrate the flexibility aggregation results of an RPG. Then, an interconnected power grid consisting of five-region interconnected power grid is used to demonstrate the coordinated dispatch results of multiple interconnected RPGs. This five-region power grid is spliced together by five European transmission systems sourced from the MATPOWER software package [30]. Detailed data are provided in the supplemental file [27]. Load curves are taken from California ISO historical data [31], and the renewable energy generation curves are obtained from the NREL website [32].
All test cases are simulated on a desktop computer with an Intel i7-10700 CPU 2.90GHz and 64 GB RAM, running on MATLAB software. The YALMIP optimization toolbox [33], MPT3 [34], solvers COPT [35] and Gurobi [36] are used to solve the optimization problems.
VI-B Flexibility Aggregation Results of Simple Case
In the IEEE 118 system, there is a DC tie-line connecting to Bus 31, and an AC tie-line connecting to Bus 17. With the flexibility aggregation method, this IEEE 118 system can be transferred into two equivalent generators with additional coupling constraints between them at each time slot.
Taking the time slot at 12:00 as an example, the coupling relationship among the boundary variables, the power of the DC tie-line connected to Bus 31, the power of the AC tie-line connected to Bus 17, and the voltage angle of Bus 17, are shown in Fig. 8. The ellipsoid represents the inner-approximated flexibility region with an ellipsoidal shape template. The gray polytope represents the actual flexibility region obtained by searching for all extreme points. The practical RPG may have far more tie-lines, making it intractable to find all the extreme points of the actual flexibility region as in this case. Therefore, the ellipsoidal shape template is a convenient tool to self-adaptively inner-approximate the flexibility region with fewer parameters.
Figure 9: Power bounds of two tie-lines in IEEE 118 RPG.
Figure 10: Ramp rate bounds of two tie-lines in IEEE 118 RPG.
In addition, the temporal coupling flexibility region across all time slots including the power and ramp rate bounds of these two tie-lines are shown in Fig. 9 and 10. The power and ramp rate limits can also be viewed as the parameters of the equivalent generators connected to these two tie-lines as illustrated in the physical counterpart in Fig. 7.
The aggregated cost function with the output power of two tie-lines as variables at 12:00 is shown in Fig. 11.
Figure 11: Aggregated cost function of IEEE 118 RPG.
VI-C Coordinated Dispatch Results of Five-Region European Power Grids
The five-region European power grid is shown in Fig. 12, with the detailed parameters listed in the supplemental file [27].
@@ VI-C Coordinated Dispatch Results of Five-Region European Power Grids
The aggregated models of these five-region European RPGs can be calculated with the proposed method. Taking the RPG R1 as an example, the power and ramp rate bounds of its seven tie-lines are shown in Fig. 13.
@@
According to Fig. 14, errors are present when using the flexibility aggregation models in non-iterative method. The errors mainly come from two aspects: the inner approximation errors between the flexibility aggregation models and the actual flexibility region, and the mismatch between the fitted cost function and the actual cost function. Despite these inaccuracies, the flexibility aggregation model-based coordination results can provide a good initial value for further iterative method’s distributed coordination calculations. This accelerates the convergence of the distributed algorithm and helps mitigate the risk of divergence in the distributed coordinated scheduling process.
VI-D Comparison with Existing Aggregation Methods
In this section, the performance of the proposed method is compared with the existing methods, including vertex search [9], multi-parametric programming [12], Fourier-Motzkin elimination [7], cubic inner approximation [13] and temporal dimension decomposition [4]. The regional power grid R1 is used as the test case. The above methods are applied to calculate the flexibility aggregation results, respectively. Furthermore, to compare the performance of the aggregated flexibility region of the RPG, the scenario coverage rate metric is used. This metric is defined as the proportion of the 100,000 operational scenarios that meet all technical constraints and are covered by the aggregated flexibility region. It serves as an indicator of the similarity between the calculated flexibility region and the actual flexibility region.
Due to the high-dimensional nature of this aggregation problem, methods like vertex search, multi-parametric programming, and Fourier-Motzkin elimination suffer from the curse of dimensionality and are unable to complete the calculation within one hour. The comparison results of other methods are shown in TABLE I.
TABLE I: Comparison of the Flexibility Aggregation Results
@@ VII Conclusion
This paper introduces a dimension-decomposition-based inner approximation method for flexibility aggregation of regional power grids with multiple AC and DC tie-lines. The aggregated model integrates the detail models of internal devices and operation constraints, facilitating optimal coordination across interconnected power grids in a non-iterative way.
To accelerate the calculation of aggregated model, enhance the coverage of the flexibility region, and derive physically meaningful flexibility aggregation results, this method incorporates three key strategies: (i) decomposing the high-dimensional flexibility aggregation region into subspaces based on the analysis of the coupling relationships among boundary variables; (ii) solving the projection problem within each subspace independently and in parallel; and (iii) employing appropriate shape templates to inner-approximate the flexibility regions.
Numerical tests demonstrate that this flexibility aggregation method is significantly less conservative than existing methods, while reliably ensuring the feasibility of the aggregated results in the dispatch process of each regional power grid.

### v3

@@ IV-B Physical Counterpart of the Flexibility Aggregation Model
The flexibility aggregation model derived through geometric projection calculations can be interpreted in terms of physical significance, as the parameters of the aggregated flexibility region of the RPG have a clear physical meaning, as shown in the schematic diagram Fig. 10.
Figure 10: (a) Schematic diagram of the original regional power grid; (b) Physical counterpart: equivalent generators model via flexibility aggregation.
For each single tie-line, its flexibility region across all time slots can be represented by an equivalent generator model, constrained by both power and ramp rate constraints, corresponding to the temporal coupling constraints in (18).
For each single time slice, the output power and bus voltage angles of all equivalent generators are interconnected through the constraints in the ellipsoidal expression form, corresponding to the single time slice coupling constraints in (21).
Therefore, after flexibility aggregation, the entire RPG can be equivalent to a group of generators that interconnected through coupling constraints.
Figure 11: Physical counterpart of IEEE 118 regional power grid: two equivalent generators model and coupling constants.
Taking the IEEE 118 RPG as an example, its physical counterpart of the flexibility aggregation model is composed of three parts: the equivalent generator of AC tie-line at Bus 17, the equivalent generator of DC tie-line at Bus 31, and the equivalent constraints among the boundary variables at each time slice, as shown in Fig. 11. Each equivalent generator can be modeled as (30). For $i\in\{17,31\}$
$\underline{p}^{\text{tie}}_{i,t}\leq p^{\text{tie}}_{i,t}\leq\overline{p}^{\text{tie}}_{i,t},\forall t\in\mathcal{T}$ (30a)
$\underline{r}^{\text{tie}}_{i,t}\leq(p^{\text{tie}}_{i,t+1}-p^{\text{tie}}_{i,t})/\Delta t\leq\overline{r}^{\text{tie}}_{i,t},\forall t\in\mathcal{T}$ (30b)
In addition, the equivalent constraints among the boundary variables at each time slice are in the ellipsoidal expression form as (31). For $t\in\mathcal{T}$
$\left\|\left(\bm{E}_{t}^{\text{bd}}\right)^{-1}\left(\bm{x}^{\text{bd}}_{t}-\bm{e}_{t}^{\text{bd}}\right)\right\|\leq 1$ (31)
where $\bm{x}^{\text{bd}}_{t}:=[p^{\text{tie}}_{17,t},p^{\text{tie}}_{31,t},\theta^{\text{bd}}_{17,t}]^{\top}$.
IV-C Calculation of the Aggregated Operational Cost Function
To facilitate coordinated economic dispatch of interconnected RPGs, the analytical aggregated cost function of RPG is derived with the piecewise linear fitting [29]. The total operational cost of an RPG is the sum of the operational costs of each generation unit. From an external system perspective, the operational cost of an RPG, denoted $\kappa_{t}$, is a function of the power flow through all tie-lines, denoted as $\bm{p}^{\text{tie}}_{t}$. To model this aggregated cost function, uniform sampling is applied in each dimension of the feasible domain, resulting in a total of $J$ samples of tie-line power flows and their corresponding operational costs. The operating point and the corresponding cost of the $j$-th sample are denoted by $\bm{p}^{\text{tie}}_{t(j)}$ and $\kappa_{t(j)}$, respectively. By solving the following optimal power flow problem, the minimum operational cost $\kappa_{t(j)}$ for the $j$-th operating point $\bm{p}^{\text{tie}}_{t(j)}$ can be obtained:
@@ IV-C Calculation of the Aggregated Operational Cost Function
In terms of the simple IEEE 118 case, the aggregated cost function with the output power of two tie-lines as variables at 12:00 is shown in Fig. 12. The aggregated cost function is a piecewise linear function expressed by the power of two tie-lines with 64 partitions.
Figure 12: Aggregated cost function of IEEE 118 RPG.
@@ V Coordinated Dispatch of Interconnected RPGs Based on the Aggregated Model
Using the aggregated flexibility models (28) and the aggregated cost function (33) of each RPG, the coordinated dispatch across multiple interconnected RPGs can be formulated as
@@ VI-A Simulation Setup
We perform simulations on a five-region European grid to demonstrate coordinated-dispatch results using our dimension-decomposition-based flexibility aggregation methods. This five-region power grid is spliced together by five European transmission systems sourced from the MATPOWER software package [31]. The five-region European power grid is shown in Fig. 13, with the detailed parameters listed in the supplementary file [28]. The load curves for simulation are taken from California ISO historical data [32], and the renewable energy generation curves are obtained from the NREL website [33].
@@
All test cases are simulated on a desktop computer with an Intel i7-10700 CPU at 2.90GHz and 64 GB RAM, running on MATLAB software. The YALMIP optimization toolbox [34], MPT3 [35], solvers COPT [36] and Gurobi [37] are used to solve the optimization problems.
VI-B Optimal Coordination Based on the Aggregated Models of RPGs
Based on the proposed methodologies, the aggregated models of these five-region European RPGs can be calculated, respectively. Taking the RPG R1 as an example, the power and ramp rate bounds of its seven tie-lines are shown in Fig. 14. These parameters can also be viewed as the parameters of the seven equivalent generators connected to these tie-lines like Fig. 10(b).
@@ VI-B Optimal Coordination Based on the Aggregated Models of RPGs
According to Fig. 15, errors are present when using the flexibility aggregation models in non-iterative approach. The errors mainly come from two aspects: the inner approximation errors between the flexibility aggregation models and the actual flexibility region, and the mismatch between the fitted cost function and the actual cost function. Despite these inaccuracies, the flexibility aggregation model-based coordination results can provide a good initial value for further iterative method’s distributed coordination calculations. This accelerates the convergence of the distributed algorithm and helps mitigate the risk of divergence in the distributed coordinated scheduling process.
VI-C Comparison with Existing Aggregation Methods
In this section, the performance of the proposed method is compared with the existing methods, including Fourier-Motzkin elimination [8], vertex search [10], multi-parametric programming [13], cubic inner approximation [14] and temporal dimension decomposition [4].
In terms of algorithmic complexity, let $T$ denote the total number of time slots considered. The Fourier-Motzkin elimination method exhibits a computational complexity of $O(T^{2^{T}})$, while both the vertex enumeration approach and multi-parametric programming have complexity on the order of $O(T^{T})$. The others exhibit exponential complexity with respect to $O(2^{T})$.
To compare the performance of the aggregated flexibility region of the RPG, we adopt the scenario coverage rate as an evaluation metric. This metric is estimated using a Monte Carlo method, which approximates the volume ratio between the inner-approximated flexibility region ^$\hat{\mathcal{R}}_{\text{L}}$ and the true feasible region $\mathcal{R}_{\text{L}}$. In this test case, it is defined as the proportion of 100,000 randomly generated operational scenarios that satisfy both technical constraints and are included within the aggregated flexibility region. The scenario coverage rate serves as an indicator of how closely the computed flexibility region approximates the actual feasible region.
Due to the high-complexity nature of this aggregation problem, methods like Fourier-Motzkin elimination, vertex search, and multi-parametric programming, suffer from the curse of dimensionality and are unable to complete the calculation within one hour. The comparison results of other methods are shown in TABLE II.
TABLE II: Comparison of the Flexibility Aggregation Results
@@ VI-C Comparison with Existing Aggregation Methods
VI-D Scalability Analysis
To evaluate the scalability of the proposed algorithm, we conducted experiments using three RPG instances with varying grid sizes and differing numbers of time slots. As shown in Fig. 16, computation time is more sensitive to the number of time slots than to the size of the grid itself. This is attributed to the fact that the number of boundary variables increases roughly in proportion to the number of time slots. These results corroborate our prior analysis, which suggests that the computational complexity increases nearly exponentially with the number of time slots.
Figure 16: Computation time across varying grid sizes and different time slots.
@@ VII Conclusion
We propose a dimension-decomposition-based inner-approximation method to aggregate the flexibility of regional grids with multiple AC/DC tie-lines. The aggregated model integrates the detail models of internal devices and operation constraints, facilitating optimal coordination across interconnected power grids in a non-iterative way.
To accelerate the calculation of the aggregated model, enhance coverage of the flexibility region, and derive physically meaningful results, we employ three key strategies: (i) decomposing the high-dimensional flexibility aggregation region into subspaces based on the analysis of the coupling relationships among boundary variables; (ii) solving the projection problem within each subspace independently and in parallel; and (iii) employing appropriate shape templates to inner-approximate the flexibility regions.
Numerical tests demonstrate that this flexibility aggregation method is significantly less conservative than existing methods while reliably ensuring the feasibility of the aggregated results in the dispatch process of each regional power grid.